OpenStudy (anonymous):

use the five-step strategy for solving word problems to find the number or numbers described in the following exercise. When a number is decreased by 75% of itself, the result is 26. What is the number? A) 78 B) 139 C) 72 D) 104

OpenStudy (anonymous):

x-75%=24 x-75/100=26

OpenStudy (anonymous):

104

OpenStudy (anonymous):

@vipul92 he said explain too.

OpenStudy (anonymous):

x*25/100=26 so x =104

OpenStudy (anonymous):

25???

OpenStudy (anonymous):

beta 100-75=25

OpenStudy (anonymous):

x-(.75x)=26 x(1-.75)=26 x(.25)=26 (x(.25))/.25=26/.25 x=26/.25

OpenStudy (anonymous):

I tried to write out all the steps.

OpenStudy (anonymous):

thank you guys!